The Talbot Company uses electrical assemblies to produce an array of small appliances. One of its high cost / high volume assemblies, the XO-01, has an estimated annual demand of 8,000 units. Talbot estimates the cost to place an order is $50, and the holding cost for each assembly is $10 per year. The company operates 250 days per year.


What is the total annual holding and ordering cost if Talbot orders using the Economic Order Quantity (EOQ)?
Answer #1

EOQ= (2*D*S/H)0.5

EOQ= (2*8000*50/10)0.5

EOQ=282.84

total annual holding and ordering cost =(D/Q)*S+(Q*H)/2

total annual holding and ordering cost =(8000/282.54)*5 + (282.84*10)/2 =$1555.77